# Digital currency hardware wallets and the essence of money

### v3

@@ 3.3 Standards and lobbying
It is plain to see why a hardware manufacturer would lobby in favour of regulations and standards that declare “offline transactions” to be an essential requirement for CBDC. But this argument is unsound: Cash is an excellent offline payment method, and to our knowledge, few if any central banks have suggested that the purpose of CBDC is to replace cash outright. The argument implicitly equivocates between a concept of “offline” that denotes transactions wherein a user is not required to carry a device with network connectivity and a concept of “offline” that involves two counterparties physically meeting in a location without third parties or network infrastructure of any kind nearby. The former scenario is the only one in which users commonly use certified hardware today. The idea that cash could suddenly become unavailable for use in the latter scenario is an overly simplistic justification for hardware manufacturers to tout their products.
@@
We note that secure elements can be used for a variety of purposes, including the prevention of exfiltration of private data from a device. However, requiring the use of secure elements for the purpose of creating attestations to parties other than the user of the device (i.e. “remote attestations”) that a device is operating within the security envelope of a trusted third party, whether a hardware manufacturer, service provider, or authority, implies a specific model for securing transactions that is not necessary for single-offline payments. For the reasons described above, we argue that single-offline payments would be both more robust and more secure if counterparties can trust network service providers, and the various methods network service providers can use to demonstrate the veracity of their claims, rather than having no choice but to rely upon trusted hardware physically held by the other counterparty. And so, for single-offline payments, trusted hardware adds nothing, but imposes significant costs, and therefore it must not be considered a requirement. Dual-offline transactions can certainly make use of remote attestation if both parties agree, but such transactions should not be seen as a general case for digital currency transactions, particularly given pervasive connectivity and the security benefits of using protocol-based security rather than introducing a requirement for counterparties to trust each other’s devices. For this reason, the development of standards that recommend or require hardware wallets to use a secure element for remote attestation are certainly problematic.
Standards that recommend hardware wallets to use a secure element to prevent the exfiltration of private data might be appropriate in the case of transactions involving smartphones, because of their complexity and large attack surface. However, it is inappropriate to assume that all or most digital wallets will be smartphones, not least because of their intrinsic security and privacy risks, despite being the “benchmark in user-accessible features”, as Barresi and Zatti argue [5].
@@ 4 Digital currency means tokens, not accounts
Barresi and Zatti assume that wallets would represent end-user money as account balances [5]. However, for ordinary consumers and non-financial businesses to hold digital currency directly, there must be some way for digital currency to be held outside accounts. Accounts generally involve a relationship between a beneficial owner (the account holder and a fiduciary, typically a regulated financial institution. A salient property of accounts is that access to assets held within accounts is determined by the identity of the account holder [28]. In contrast, tokens are assets that can be held outside of accounts. Tokens can be bearer instruments, meaning that they can be used in a way that does not index the identity of their owners. Payments can be held offline and transacted online, so the user experience would be similar to an EMV card payment except in which the payer presents digital assets rather than instructions to debit an account.
@@
Compared to assets held within accounts, assets that are held directly by their owners have a different set of security requirements for their owners. For example, owners of assets held within accounts can rely upon various identification mechanisms, such as PIN numbers and two-factor authentication, to protect their assets if their cards or mobile devices are stolen, and asset custodians can also provide credit to account holders, such that if assets are spent unlawfully, the account holders can ask the asset custodians cover the value of their losses. In contrast, the bearer is responsible for bearer instruments. If cash is lost or stolen, owners generally have little recourse beyond law enforcement mechanisms to recover their money, and such mechanisms often fail to recover the value. The potential for loss or theft exists not only physically but also electronically. Devices can malfunction, and devices can be compromised. The prevailing model for handling devices that no longer work or have been compromised is to wipe their contents completely or to replace them wholesale. When such devices are used to access accounts, such replacement is not much of a problem, although bearer instruments inexorably expose their owners to new risks that require different solutions.
Some central banks, including the ECB [14], the US Federal Reserve [29], and the Bank of England [31], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs. Such arguments are nonsense. Limiting CBDC withdrawals, rather than aggregate holdings, would be no less effective in preventing bank runs, and the approach of limiting bank withdrawals has been used successfully in the past, notably in Greece in 2015 [30]. Importantly, limiting CBDC withdrawals can be done without requiring wallets to implement account-like control mechanisms that undermine the will of their users. Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [5]; little could be more disrespectful of a user’s autonomy.
Although a third party is necessary for fair exchange (and indeed trusted hardware or closed-source software blobs can be this third party de facto), Pagnia and Gärtner have shown that the role of the third party can be limited, even zero-knowledge: tokens can be transferred without the third party knowing what is being transferred or having a custodial relationship with either party [20].
@@
The idea that digitalisation implies the use of accounts is not true in general. Digital tokens can be real indeed, with ownership demonstrable by transferring it: using a single-use private key to transfer the asset to another single-use private key. The requirement of an account in the form of a “wallet provider” is not true for cryptocurrencies in general, nor does a private key constitute a proof of identity when it is created to be used exactly once. It is true that many Bitcoin users rely on service providers to manage their holdings. However, to argue from this that self-custody of tokens will never take off is pure speculation, and it ignores the fact that most Bitcoin users are speculators with no expectation that they would use Bitcoin for ordinary retail purposes. Many people carry around an amount of cash that is useful for retail purchases but small enough that they could lose it without too much regret, despite the fact that losing it is forever.
Digital assets can certainly be held by users without accounts of any kind, for example, as a payer can in Chaum’s model. Accounts for recipients might be desirable, but they are not technically necessary. The involvement of a third-party to prevent double-spending does not imply that the third party must have a role as custodian or account provider, or knowledge of the asset being transferred, the circumstances of the transfer, or the identity of either party.
@@
Tokens are also important for privacy. Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [36, 37], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules. This means that transactions are not independent of one another. The property of transaction independence is important because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [38].
@@
Finally, the idea that digital currency intrinsically requires accounts is not true; central banks can certainly issue tokens that exist outside of accounts, just as they do with cash, and a token like this would be a liability on a central bank’s balance sheet, not unlike cash or central bank reserves.
@@ 5 Separating issuance from transaction processing
Even if transactions are done with network access to a mutually trusted third party, there is no reason to assume that the issuer must be involved in transaction processing in real-time. Chaum’s original proposal assumed such a dependency [33], but this is not a requirement in general. In particular, it is possible to register a (zero-knowledge) commitment to transfer an asset without communicating with the issuer, and for a transaction to have finality it is sufficient for both parties have faith in a mutually trusted third party to eventually post the commitment to a bulletin board recognised by the issuer [10]. Also, if tokens carry their own provenance information, then there is no need for the issuer to track tokens that it had seen, either [10].
It is certainly possible to transact money issued by a central bank in a decentralised system, and wallets can be expected to respect this possibility. Trusting a central bank (or other issuing authority, such as those that issue e-money) to issue money and maintain its integrity is certainly different from trusting it to process transactions. In particular, it is possible to use a decentralised system to transact centrally issued money. The use of a secure element to establish trust implicitly requires the transacting parties to trust a hardware manufacturer or other central authority to facilitate transactions. Such a role for a unified central authority might be incompatible with the concept of a decentralised system. Finally, requirements for hardware wallets used for digital currency should not be needlessly inconsistent with requirements for hardware wallets used for other digital assets.
@@
It is also worth considering that all users cannot be expected to trust the same trust anchors equally. Although the fact that assets held in offline non-custodial wallets can be transacted online is sufficient to reject the requirement of a hardware root of trust for security, even the case of dual-offline transactions poses a problem for requiring the use of secure elements. Specifically, there is the matter of who the transacting parties choose to trust. We imagine that parties choosing to transact online could potentially choose from a variety of manufacturers of hardware roots of trust based upon their needs and security models, whereas forcing manufacturers of hardware wallets to embed hardware certified by a specific authority, with the expectation that this authority is universally and equally trusted by all users, is not appropriate. We believe that allowing users to retain a choice of mutually trusted hardware security solutions, such as removable eSIM devices, is appropriate, although it might be impractical for users to carry around as many hardware devices as there are commonly trusted third parties with their various counterparties. The fact that hardware devices impose a trust model based upon their specific manufacturer is yet another reason to reject the idea that such devices could function as a public utility.

### v4

@@ 3.3 Standards and lobbying
It is plain to see why a hardware manufacturer would lobby in favour of regulations and standards that declare “offline transactions” to be an essential requirement for CBDC. But this argument is unsound: Cash is an excellent offline payment method, and to our knowledge, few if any central banks have suggested that the purpose of CBDC is to replace cash outright. The argument implicitly equivocates between a concept of “offline” that denotes a transaction wherein a user is not required to carry a device with network connectivity, versus a concept of “offline” that denotes a transaction between two counterparties that physically meet in a location without third parties or functioning network infrastructure of any kind nearby. The former scenario is the only one in which users commonly use certified hardware today, for example, single-offline POS payments (e.g. Apple Pay) and network applications that rely upon DRM (e.g. video streaming). Even the specialised Suica card offered by Japan Rail for use at its payment gates, which uses embedded trusted hardware to facilitate payments when connectivity to a central server is broken [Shiibashi 2008], could potentially use locally reachable points of trust to aggregate transaction information rather than assume that network connectivity is either all or nothing. Moreover, the idea that cash could suddenly become unavailable for general use in offline situations (the latter scenario) is an overly simplistic justification for a system-level requirement.
@@
We note that secure elements can be used for a variety of purposes, including the prevention of exfiltration of private data from a device. However, requiring the use of secure elements for the purpose of creating attestations to parties other than the user of the device (i.e. “remote attestations”) that a device is operating within the security envelope of a trusted third party, whether a hardware manufacturer, service provider, or authority, implies a specific model for securing transactions that is not necessary for single-offline payments. For the reasons described above, we argue that single-offline payments would be both more robust and more secure if their locus of trust can be a network service, relying upon the various methods network service providers can use to demonstrate the veracity of their claims, rather than having no choice but to rely upon the integrity of trusted hardware physically held by the other counterparty.
locus of trust network service certified hardware
supports secure single-offline payments yes yes
supports secure dual-offline payments no yes
requires wallets to be linked to attributed accounts no yes
allows anonymous payers (bearer instruments) yes no
allows unregistered or uncertified end-user devices yes no
devices may be built by user communities without approval yes no
devices may be fully auditable by end-users yes no
Table 2: Features of a network-based trust locus versus a hardware-based trust locus.
Table 2 summarises the key differences between using a network service as a root of trust and using certified hardware as root of trust. Although certified hardware has the advantage of enabling dual-offline payments, such payments are not actually common in the digital economy. Secure network services can provide the same degree of transaction assurance that secure hardware devices can provide, with the only limitation being that they must be reachable by one of the transacting parties. At the same time, we have argued that because it is feasible for adversarial state actors or organised crime syndicates to compromise individual devices in their possession, it would be too much of a risk for a system relying upon such devices to allow them to spend money that is not linked to attributed accounts, because compromised devices would be able to reuse anonymous money (bearer instruments) without limit. However, with a network service, there is no system-level risk if an end-user device is compromised, since the locus of trust is elsewhere. A network service that provides payment finality222the state in which obligations are discharged between transacting parties. for bearer instruments can be operated by a trusted third party, as in GNU Taler, or it can be oblivious and based on public commitments, so that compromising it does not enable double-spending [Friolo 2025].
And so, for single-offline payments, trusted hardware adds nothing, but imposes significant costs and limitations, and therefore it must not be considered a system-level requirement. Dual-offline transactions can certainly make use of remote attestation if both parties agree, but such transactions should not be seen as a general case for digital currency transactions, particularly given pervasive connectivity and the security benefits of using protocol-based security with an external root of trust rather than introducing a requirement for counterparties to trust each other’s devices. For this reason, the development of standards that recommend or require hardware wallets to use a secure element for remote attestation are certainly problematic, while perfectly functional wallet devices that provide better end-user affordances could be designed to strictly facilitate single-offline payments instead. Such devices do not necessarily require network capabilities themselves, provided that their transaction counterparties are online.
Standards that recommend hardware wallets to use a secure element to prevent the exfiltration of private data might be appropriate in the case of transactions involving smartphones, because of their complexity and exposure to malware attacks. However, it is not necessarily the case that all or most digital wallets will be smartphones, despite being the “benchmark in user-accessible features”, as Barresi and Zatti have argued [Barresi 2023]. One reason is price: Although virtually everyone must have a way to engage with the economy, a substantial share of the population cannot afford a smartphone. In 2024, the global median price of a smartphone was US\$92.59, representing 11% of global median per capita income [GDIP 2024]. Another reason is security: with ever more complex operating systems and applications, smartphones are increasingly plagued by security vulnerabilities [NIST], many of which specifically target banking credentials [Fadilpašić 2025]. For a sensitive function like payments, it might be safer and more appropriate to use a dedicated device [Bizama 2025].
@@ 4 Digital currency means tokens, not accounts
Barresi and Zatti assume that wallets would represent end-user money as account balances [Barresi 2023]. However, for ordinary consumers and non-financial businesses to hold digital currency directly, there must be some way for end-users to hold digital currency without the requirement to have an attributed account. Accounts generally involve a relationship between an account holder and a fiduciary, typically a regulated financial institution. A salient property of accounts is that access to assets held within accounts is determined by the identity of the account holder [Auer and Böhme 2020], and the balance of an account represents an entitlement that can be adjudicated, for example, in a court of law. In contrast, tokens are assets that can be held outside of accounts. Tokens can be bearer instruments, meaning that they can be used in a way that does not reference the identity of their holders. Payments can be held offline and transacted online, so the user experience would be similar to an EMV card payment except in which the payer presents digital assets rather than instructions to debit an account.
@@
Compared to assets held within accounts, assets that are held directly by their holders have a different set of security requirements. For example, holders of assets held within accounts can rely upon various identification mechanisms, such as PIN numbers and two-factor authentication, to protect their assets if their cards or mobile devices are stolen, and asset custodians can also provide assurances to account holders that if assets are spent against their wishes, then the account holders can ask their asset custodians cover the value of their losses. In contrast, the bearer is responsible for bearer instruments. Persons whose cash is lost or stolen generally have little recourse beyond law enforcement mechanisms to recover their money, and such mechanisms often fail to recover the value. The potential for loss or theft exists not only physically but also electronically. Devices can malfunction, and devices can be compromised. The prevailing model for handling devices that no longer work or have been compromised is to wipe their contents completely or to replace them wholesale. When such devices are associated with accounts, replacing devices that have been demonstrated to be insecure might be operationally expensive or cumbersome, but the actions of a compromised device are ultimately traceable to the account with which it is associated, thus giving the account holder ways to mitigate the effects. Although the potential for loss or theft of bearer instruments inexorably constitutes a risk that requires a different kind of solution, approaches well-suited to digital technology, such as backup copies and withdrawal records that can be used by victims of theft to selectively de-anonymise their own tokens that had been stolen, can be robust and effective in mitigating such risks.
Some central banks, including the ECB [European Commission 2023], the US Federal Reserve [Federal Reserve 2022], and the Bank of England [Bank of England 2023], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs. However, as an alternative to the heavy-handed approach of monitoring or restricting wallets, an approach of limiting CBDC withdrawals from banks would be no less effective in preventing bank runs or bank disintermediation than limiting cash withdrawals from banks, and the approach of limiting bank withdrawals from banks has been used successfully in the past, notably in Greece in 2015 [Associated Press 2018]. In this sense, CBDC is no different from cash. To argue that requiring wallets to be managed by parties other than their holders is necessary to address financial stability problems resulting from reduced liquidity is to conflate the technical design requirements for a wallet with economic justifications for eliminating the ability for individuals to hold money outside accounts. However, although some banks and governments might see value in eliminating cash-like instruments and might view CBDC as a vehicle for achieving that objective, such an argument is not commonly articulated.
Importantly, limiting CBDC withdrawals can be done without requiring wallets to implement account-like control mechanisms that undermine the will of their users. Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [Barresi 2023]; little could be more disrespectful of a user’s autonomy. Similarly, the argument that digital currency could be used as a way to implement “programmable money” that can only be spent in pre-defined ways is a non-starter. Such a mechanism would not only undermine the singleness of money but also open a Pandora’s box of ethical concerns [Li 2025], some of which have already been realised by the failed cashless debit card scheme in Australia [Greenacre 2023]. However, banks could impose simple rules to limit the aggregate size of withdrawals that can be made from their accounts into CBDC, both in a single transaction and over a period of time, and such rules could be potentially combined with identity information to limit the amount that an end-user could withdraw throughout the system. Such rules could also be combined with periodic token expiry, as a mechanism to prevent hoarding [Goodell 2021]. Periodic token expiry is less intrusive than mechanisms that restrict or control wallets directly, and it can be made transparent to users by implementing a mechanism for wallets to automatically refresh any assets that are close to the expiration time.
Although a third party is necessary for fair exchange (and indeed trusted hardware or closed-source software blobs can act as this third party de facto), Pagnia and Gärtner have demonstrated that the role of the third party can be performed “in zero-knowledge”. Specifically, a third party can facilitate final settlement of an asset transfer involving two transacting parties, without having a custodial relationship with either party, without knowing the identity of either party, without knowing how much is being transferred, and even without knowing that tokens are being transferred at all [Pagnia 1999]. Importantly, with oblivious management, an issuer does not rely upon a central ledger to track tokens individually, although the issuer must trust that the system that collects transactions and produces rolled-up aggregation entries, one per time period, will not equivocate. This trust can be enhanced by requiring that system to externalise its commitment to a single history, for example by sending some or all of its aggregation entries to the issuer, or by writing them to a public bulletin board or distributed ledger.
@@
The idea that digitalisation implies the use of accounts is not true in general. Digital tokens can indeed be held directly as data, with ownership demonstrable by transferring it: using a single-use private key to transfer the asset to another single-use private key. A transaction is a cryptographically signed mapping of that single-use key to a commitment to assign the asset to another single-use key, and the “proof of provenance” offered by an integrity provider demonstrates that the first single-use key has been irreversibly mapped in this way. Users will require a mechanism for generating keys, but open source tools plus a good source of entropy are sufficient for this. The requirement of an account in the form of a “wallet provider” is not true for cryptocurrencies in general, nor does a private key constitute a proof of identity when it is created to be used exactly once. It is true that many Bitcoin users rely on service providers to manage their holdings. However, to argue from this that self-custody of tokens will never take off is pure speculation. Many people deliberately carry around an amount of cash that is useful for retail purchases but small enough that they could lose it without too much regret, despite the fact that losing it is forever. They exercise a free choice to determine how much cash to carry, without being forced via technology to adhere to some limit. More importantly, even those who might be comfortable with a limit might be uncomfortable with the mechanism used to impose the limit, particularly if it requires them to use a device controlled by someone else (as described in Section 3.2).
Digital assets can indeed be held by users without accounts of any kind, just like cash and other physical assets. An example of a digital asset that can be held without an account is an e-cash token in Chaum’s model [Chaum 1982, Chaum 2021], which requires every transaction to function as a redemption, and it has also been shown that digital assets can be transferred without redemption [Goodell 2023, Friolo 2025]. Accounts for recipients might be desirable, but they are not technically necessary. As described above, it has been proven that the prevention of double-spending does not imply the existence of a traceable record of all the holdings and movements through sequential time. In Chaum’s original proposal, an asset can be transferred to a recipient without revealing information about the payer, and, unlike oblivious management, double-spending is prevented by maintaining a record of the spent tokens. In the general case, and in the approach taken by oblivious management, to prevent double-spending, it is sufficient to register non-repudiable commitments from payers, which may take the form of a hash or zero-knowledge proof that can be verified without knowledge of the transaction [Pagnia 1999], and proofs of these commitments can be delivered along with the asset data rather than centrally recorded [Goodell 2023, Friolo 2025]. Oblivious management presupposes that issuers can be held responsible for what they have issued and that proofs of commitments that cryptographically demonstrate the validity and integrity of an asset constitute sufficient evidence of an issuer’s responsibility to honour a redemption. As discussed earlier, specific payment architectures that implement these mechanisms have been proposed.
@@
Tokens are also important for privacy. Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [Wüst 2022, Kiayias 2022], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules. This means that transactions are not independent of one another. The property of transaction independence is important, because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [Friolo 2025]. Because it must be possible for balances to be tracked through all of the movements of that balance, transactions in systems that represent money as balances cannot achieve transaction independence. However, systems that implement money as tokens are not subject to this constraint.
@@
Finally, the idea that digital currency intrinsically requires accounts is not true; central banks can certainly issue tokens that exist outside of accounts, just as they do with cash, and a token like this would be a liability on a central bank’s balance sheet, not unlike cash.
@@ 5 Separating issuance from transaction processing
Even if transactions are done with network access to a mutually trusted third party, there is no reason to assume (a) that the third party must be the issuer, or (b) that the issuer must be involved in transaction processing in real-time. Chaum’s original proposal assumed both conditions [Chaum 1982], but they are not requirements in general. In particular, with oblivious management, it is possible to register a (zero-knowledge) commitment to transfer a digital asset without communicating directly with the issuer. For a transaction to have finality, assuming every transfer specifies the payee with a unique single-use key, it is sufficient for both parties have confidence in a mutually trusted third party, acting as a notary, to eventually post the commitment (or an aggregation that comprehends the commitment) to an append-only bulletin board that maps single-use keys to commitments and that is trusted by the issuer [Goodell 2023]. The append-only bulletin board can be operated by the issuer directly, by a party that the issuer trusts, or as a distributed ledger operated by a set of parties that the issuer trusts in aggregate. Provided that the operator of the bulletin board does not equivocate, then since no two tokens can be transferred to the same single-use key and have a valid proof of provenance, it is not necessary for the issuer to maintain a ledger that records the registration status of individual tokens [Goodell 2023].
Even if we assume that all tokens are issued by a central issuer, such as a central bank, it is certainly possible to use a decentralised system to conduct transactions using those tokens, and wallets can be designed to respect this possibility. Trusting a central bank (or other issuing authority, such as one that issues e-money) to issue currency and maintain its integrity is certainly different from trusting it to process e-cash transactions. In particular, it is possible to use a decentralised system to transact centrally issued money. Just as cash can be managed with a two-tiered approach, e-cash can also be managed with a two-tiered approach. If we assume, following Chaum, that individual e-cash assets are generated for one-time use by consumers, then a central bank can provide private-sector banks with “genesis” assets that can be burned, pari passu, by private-sector banks in the process of creating single-use e-cash tokens for consumers. Then, the single-use assets can be recirculated within the system by private-sector banks indefinitely via the same process before ultimately being redeemed with the central bank [Friolo 2025]. The use of a secure element to assure trust seems to implicitly require the transacting parties to respect the authority of a specific hardware manufacturer, and the power implicit to this authority might be incompatible with the concept of decentralisation if all transacting parties are forced to use secure elements produced by a limited set of manufacturers. To require the use of a hardware root of trust with remote attestation unnecessarily limits decentralisation.
The fact that assets held in offline non-custodial wallets can be transacted online is sufficient to reject the requirement of a hardware root of trust for security, because trusted hardware is unnecessary for online and single-offline payments (see Table 2), and not every transaction must be treated as a dual-offline transaction. However, even if we accept that trusted hardware is necessary for dual-offline payments, a key problem remains: Which hardware manufacturer should transacting parties choose to trust? The risk implicit to requiring all users of a system to respect the same set of trusted authorities is well-documented [Kushner 2013]. The incentives for operators of infrastructure, such as telecommunications providers operating under public monopoly, to discourage beneficial innovation are also well-documented [Tran 2019].
@@
We imagine that parties choosing to transact online could potentially choose from a variety of trust providers based upon their needs and security models, whereas forcing manufacturers of hardware wallets to embed hardware certified by a specific authority, with the expectation that this authority is universally and equally trusted by all users, may be too demanding. We believe that allowing users who need dual-offline payments to retain a choice of mutually trusted hardware security solutions, such as removable eSIM devices, is appropriate, although it might be impractical for users to carry around as many hardware devices as there are commonly trusted third parties with their various counterparties. Interoperability does not solve this problem, since it is the hardware itself that must be trusted. Substituting consumer choice with public infrastructure might seem like a tempting solution, although it entails requiring consumers to relinquish some of their autonomy.
Finally, there is no reason for which the technical requirements for hardware wallets used for digital currency in general should be inconsistent with the technical requirements for hardware wallets used for other transferable digital assets, including those that do not require a hardware root of trust because they are not expected to support dual-offline payments. Both kinds of wallets, those that satisfy special requirements to support dual-offline payments and those that do not, should be able to work within a digital currency system.